OpenStudy (anonymous):

The perimeter of a face of a cube is 16 mm. What is its volume?

OpenStudy (anonymous):

face of cube is in shape of square

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

so perimeter of a square = 4a

OpenStudy (anonymous):

okey

OpenStudy (anonymous):

4a=16

OpenStudy (nurali):

The perimeter of the face is the same as the perimeter of a square: you add up the lengths of each side. There are four equal sides that add up to 16. 4S=16 (as an equation) or if you like S+S+S+S=16 either way, S=4 the volume of a cube is S 3 or S*S*S. So volume is 4*4*4= 16*4= 64 mm 3

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

thanks guys

OpenStudy (anonymous):

there u go